# supplement

### v1

@@ I Introduction
In this work, we present a differentiable implementation of the local natural orbital coupled cluster (LNO-CC) method, first introduced by Rolik and Kállay,[17] within the AD framework of PySCFAD. The modular design of PySCFAD allows for the integration of new methods with virtually no changes to the existing components of the software. Furthermore, performance optimizations can be targeted to a small segment of the program, e.g., the tensor contractions for computing the triple excitation correction in the CCSD(T) method, without compromising the differentiability of the entire computational workflow. We demonstrate the resulting differentiable LNO-CC method in some non-trivial applications: calculating the bond orders and Mössbauer spectroscopy parameters for a model system of the [NiFe]-hydrogenase active site, as well as obtaining the anharmonic infrared (IR) spectra of a protonated water cluster from ab initio molecular dynamics.
@@ II Notations
Throughout the paper, we use Greek letters ($\mu$, $\nu$, …) to denote atomic orbitals. Occupied and virtual canonical Hartree-Fock (HF) orbitals are labeled using lowercase letters ($i$, $j$, … for occupied; $a$, $b$, … for virtual), while uppercase letters ($I$, $J$, …) signify local orbitals. The energies corresponding to canonical HF orbitals and semi-canonical local orbitals are denoted by $\varepsilon$ and $\mathcal{E}$, respectively. Orbitals that span a local active space are distinguished by tildes ($\tilde{i}$, $\tilde{j}$, …). Unless specified otherwise, equations are expressed in the spin orbital formalism, although our implementation is spin-adapted, and we adopt Dirac’s notation for the electron integrals.
@@ IV.1 Nuclear Gradient and Dipole Moment
It can be seen that, on average, the errors in nuclear gradients and dipole moments are of the same order as those in energies. We note that both nuclear gradients and dipole moments are types of first-order response properties, which can be determined using the zeroth-order wavefunction (or density matrix), according to Wigner’s $12n+$ 1 rule. However, for higher order response properties, such as polarizabilities, which depend on the first-order wavefunction, it may be necessary to use larger correlation domains to achieve the desired accuracy, as suggested by Crawford and co-workers.[65]
@@ IV.2 Geometry Optimization
The present implementation of LNO-CCSD(T) has a formal scaling of $O(N^{5}+\tilde{N}^{7})$, where $N$ is the number of basis functions, and $\tilde{N}$ represents the dimensions of the local correlation domains. The $O(N^{5})$ part of the computational complexity originates from the global MP2 correction and the LNO construction scheme, while the $O(\tilde{N}^{7})$ part is due to the local CCSD(T) calculations. For the chemical systems for which one would use CCSD(T), $\tilde{N}$ does not need to grow with the system size for a given energy accuracy per atom, making LNO-CCSD(T) orders of magnitude more efficient than canonical CCSD(T), especially for large $N$. However, employing a looser LNO cutoff (and thus a smaller $\tilde{N}$) might introduce more severe discontinuities into the potential energy surface, leading to more iterations to converge the geometry optimization. Overall, our tests suggest that our LNO-CCSD(T) method is capable of accurate geometry optimizations with computational scaling comparable to that of canonical MP2, albeit potentially with a large prefactor.
@@ IV.3 Relaxed Density Matrix
In Table 2, the orbital resolved bond orders (defined in Eq. LABEL:eq:bond_order) are presented for the two metal centers. Minor contributions to bonding are omitted, and only representative bonds among those with similar chemical environments are shown. We compare the results computed by DFT with those obtained by the LNO-CCSD(T) method. For the latter, IAOs were employed as the local orbitals, and an LNO cutoff of $\zeta=2\times 10^{-5}$ was adopted, leading to fragments with correlation domains comprising approximately $80$ occupied and $200$ virtual orbitals at most. (Such choices are discussed in the supporting information Sec. LABEL:sec:HtSH.) The key findings are the following: (i) Bond orders calculated by the two methods qualitatively agree with each other. (ii) The $4s$ orbitals of both Fe and Ni do not contribute significantly to bonding. (iii) Relatively strong covalent bonding is observed between Fe and CO. (iv) The three bridging cysteines exhibit weak bonding interactions with the two metal centers, characterized by the stronger Ni–S bonds compared to the Fe–S bonds. Interestingly, the LNO-CCSD(T) method tends to predict non-bonding characters for the Fe–S bonds.

### v3

@@ I Introduction
In this work, we present a differentiable implementation of the local natural orbital coupled cluster (LNO-CC) method, first introduced by Rolik and Kállay,[17] within the AD framework of PySCFAD. The modular design of PySCFAD allows for the integration of new methods with virtually no changes to the existing components of the software. Furthermore, performance optimizations can be targeted to a small segment of the program, e.g., the tensor contractions for computing the triple excitation correction in the CCSD(T) method, without compromising the differentiability of the entire computational workflow. We demonstrate the resulting differentiable LNO-CC method in some non-trivial applications: calculating the bond orders and Mössbauer spectroscopy parameters for a model system of the [NiFe]-hydrogenase active site, as well as obtaining the anharmonic infrared (IR) spectra of a protonated water hexamer from ab initio molecular dynamics.
@@ II Notations
Throughout the paper, we use Greek letters ($\mu$, $\nu$, …) to denote atomic orbitals. Occupied and virtual canonical Hartree-Fock (HF) orbitals are labeled using lowercase letters ($i$, $j$, … for occupied; $a$, $b$, … for virtual), while uppercase letters ($I$, $J$, …) signify local orbitals, whether semi-canonicalized or not. The energies corresponding to canonical HF orbitals and semi-canonical local orbitals are denoted by $\varepsilon$ and $\mathcal{E}$, respectively. Orbitals that span a local active space are distinguished by tildes ($\tilde{i}$, $\tilde{j}$, …). Unless specified otherwise, equations are expressed in the spin orbital formalism, although our implementation is spin-adapted, and we adopt Dirac’s notation for the electron integrals.
@@ IV.1 Nuclear Gradient and Dipole Moment
It can be seen that, on average, the errors in nuclear gradients and dipole moments are of the same order as those in energies. Enlarging the basis set from double zeta to triple zeta does not lead to an increase in the errors (as shown in Fig. LABEL:fig:grad_pm_TZ). We note that both nuclear gradients and dipole moments are types of first-order response properties, which can be determined using the zeroth-order wavefunction (or density matrix), according to Wigner’s $12n+$ 1 rule. However, for higher order response properties, such as polarizabilities, which depend on the first-order wavefunction, it may be necessary to use larger correlation domains to achieve the desired accuracy, as suggested by Crawford and co-workers.[65]
@@ IV.2 Geometry Optimization
The present implementation of LNO-CCSD(T) has a formal scaling of $O(N^{5}+\tilde{N}^{7})$, where $N$ is the number of basis functions, and $\tilde{N}$ represents the dimensions of the local correlation domains. The $O(N^{5})$ part of the computational complexity originates from the global MP2 correction and the LNO construction scheme, while the $O(\tilde{N}^{7})$ part is due to the local CCSD(T) calculations. For the chemical systems for which one would use CCSD(T), $\tilde{N}$ does not need to grow with the system size for a given energy accuracy per atom, making LNO-CCSD(T) orders of magnitude more efficient than canonical CCSD(T), especially for large $N$ (see Fig. LABEL:fig:time_pm_TZ and the corresponding discussion). However, employing a looser LNO cutoff (and thus a smaller $\tilde{N}$) might introduce more severe discontinuities into the potential energy surface, leading to more iterations to converge the geometry optimization. Overall, our tests suggest that our LNO-CCSD(T) method is capable of accurate geometry optimizations with computational scaling comparable to that of canonical MP2, albeit potentially with a large prefactor.
@@
a The geometry used for benzene was obtained from iteration 100 of the geometry optimization.
@@ IV.3 Relaxed Density Matrix
In Table 2, the orbital resolved bond orders (defined in Eq. LABEL:eq:bond_order) are presented for the two metal centers. Minor contributions to bonding are omitted, and only representative bonds among those with similar chemical environments are shown. (See Tables LABEL:tab:NiFe_bond_order1_full and LABEL:tab:NiFe_bond_order for the complete data.) We compare the results computed by DFT with those obtained by the LNO-CCSD(T) method. For the latter, IAOs were employed as the local orbitals, and an LNO cutoff of $\zeta=2\times 10^{-5}$ was adopted, leading to fragments with correlation domains comprising approximately $80$ occupied and $200$ virtual orbitals at most. (Such choices are discussed in the supporting information Sec. LABEL:sec:HtSH.) The key findings are the following: (i) Bond orders calculated by the two methods qualitatively agree with each other. (ii) The $4s$ orbitals of both Fe and Ni do not contribute significantly to bonding. (iii) Relatively strong covalent bonding is observed between Fe and CO. (iv) The three bridging cysteines exhibit weak bonding interactions with the two metal centers, characterized by the stronger Ni–S bonds compared to the Fe–S bonds. Interestingly, the LNO-CCSD(T) method tends to predict non-bonding characters for the Fe–S bonds.